is there any other multiplication problem that equals the same as 8 * 1274?
Found 2 solutions by checkley77, edjones:
Answer by checkley77(12844)

8 x 1274=10,192
16*637=10,192
112*91=10,192

Answer by edjones(8007)

Lots!
8*1274
=4*2*1274
=4*2548
=2*5096
=1*10192
=16*637
.
Ed